a 5 cm^3 solution of h2o2 liberates 0.508g of iodine from an acidified KI solution. what is the volume strength of h2o2 at stp?
 

Grade:10

1 Answers

Arun
6 years ago
The redox reaction involved is : H2O2 + 2I- + 2H+ → 2H2O + I2 If M is molarity of H2O2 solution, then 5M = (0.508 × 1000)/254 (∵ 1 mole H2O2 ≡ 1 mole I2) ⇒ M = 0.4 Also, n-factor of H2O2 is 2, therefore normality of H2O2 solution is 0.8 N. ⇒ Volume strength = Normality × 5.6 = 0.8 × 5.6 = 4.48 V